OpenStudy (clamin):

ONE MEDAL!!please Identify whether the question is a permutation or combination..then solve the problem please..How many baseball teams of 9 members can be formed from 14 payers?

OpenStudy (kropot72):

The order in which the players are selected doesn't matter. Therefore we require the number of combinations of 14 players selected 9 at a time: \[14C9=\frac{14!}{5!9!}=\frac{14\times13\times12\times11\times10}{5\times4\times3\times2}=you\ can\ calculate\]

OpenStudy (clamin):

so is it okay tp calculate 14 through 10 not including 9 til 1??

OpenStudy (kropot72):

In simplifying the fraction with factorials, 14! divided by 9! gives 14 * 13 * 12 * 11 * 10. Then the denominator becomes 5! = 5 * 4 * 3 * 2.